OpenStudy (anonymous):

use the descriminant to determine how many real roots each has. 4x^2-28x+49=0

OpenStudy (lgbasallote):

1

OpenStudy (anonymous):

can you show me how you got your answer

OpenStudy (anonymous):

a miracle occurs and when you compute \[b^2-4ac=(-28)^2-4\times 4\times 49=0\]meaning a) one real root and b) you have a perfect square

OpenStudy (anonymous):

it is \[(2x-7)^2\]

OpenStudy (anonymous):

thank you